1) Assume that the cost of a unit of labor (W) is $50 (the nominal wage) and a unit of labor 

produces one unit. Calculate the price per unit if mark up is :


i. 20%


ii. 40%


1
Expert's answer
2022-01-14T09:19:40-0500

"P_1=\frac{50}{1+\frac{20}{100}}=41.7" $.

"P_2=\frac{50}{1+\frac{40}{100}}=35.7" $.